OpenStudy (anonymous):

(1 - cos^2x)/(1 + cot^2x) = sin^2x true or false?

OpenStudy (raden):

hint : 1-cos^2 x = sin^2 x 1 + cot^2x = csc^2 x

hero (hero):

(1 - cos^2x)/(1 + cot^2x) = sin^2x 1 - cos^2x = sin^2x(1 + cot^2x) 1 - cos^2x = sin^2x + sin^2xcot^2x 1 - cos^2x = sin^2x + cos^2x 1 - cos^2x = 1 1 - 1 = cos^2x 0 = cos^2x False

hero (hero):

However, it is true whenever x = pi/2

OpenStudy (raden):

i got for left side be sin^4 x so, LHS not equal RHS

hero (hero):

I already stated that it was false

hero (hero):

There's always another way

OpenStudy (raden):

yeps.. :)

hero (hero):

I hope I never have to teach trig because I don't like doing these the other way.

OpenStudy (anonymous):